# CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery

### v1

@@ IV-A Qualitative and Quantitative Analysis
TABLE II: Performance comparison of image enhancement methods for various computer vision tasks.
Method ORB Matches (Lowe’s Test) Object Detection (Visible Objects) ArUco Marker Detection
Correct Incorrect Detected Missed Mislabelled Correct Missed
Retinex Theory [27] 49 451 1 9 0 5 1
CLAHE [28] 36 464 1 9 0 3 3
Zero-DCE [29] 61 439 0 10 0 5 1
LLFormer [30] 53 447 0 10 0 5 1
CIDNet [31] 66 434 1 9 0 6 0
Raw IR 31 469 0 9 1 0 6
CLEAR-IR (Proposed Method) 75 342 7 2 1 6 0
IV-B Feature Matching and Keypoint Detection
To assess the impact of emitter pattern removal on traditional computer vision algorithms, the performance of the Oriented FAST and Rotated BRIEF (ORB) [32] feature detector was evaluated. ORB is a robust algorithm for finding keypoints and feature descriptors, which are fundamental for tasks such as Simultaneous Localisation and Mapping (SLAM) and visual odometry.
As shown in Figure 6, the raw IR images exhibit numerous false keypoint detections due to the high-frequency nature of the emitter pattern. These spurious features lead to inaccurate and unreliable feature matches, which would compromise downstream tasks. In contrast, the denoised images from the CLEAR-IR model provide a clean representation where ORB is able to detect and match meaningful keypoints on actual scene structures, such as object boundaries and corners. This demonstrates that the proposed method effectively suppresses noise that could be misinterpreted as texture by feature detectors, thereby producing a more reliable input for foundational vision algorithms.
(a) ORB matches on raw IR image
(b) ORB matches on denoised IR image
Figure 6: ORB feature matching on (a) raw IR and (b) denoised IR images.
IV-C Object Detection and Segmentation
@@
To demonstrate the utility of the denoised images for advanced perception tasks, a pre-trained YOLOv8 object detection and segmentation model was applied. This model was not trained on IR data, making this a zero-shot evaluation to determine if the proposed preprocessing step can make the images usable with standard RGB-based vision pipelines.
(a) YOLOv8 on RGB image
(b) YOLOv8 on raw IR image
(c) YOLOv8 on Retinex image
(d) YOLOv8 on CLEAR-IR image
@@
Figure 7: Object detection and segmentation results of YOLOv8 on raw and denoised IR images.
@@
On the raw IR images, YOLOv8 failed to detect or segment objects. The emitter patterns and resulting image noise prevented the model from recognising the distinct shapes and features of common objects. After preprocessing the images with the CLEAR-IR model, the performance of YOLOv8 dramatically improved. As illustrated in Figure 7, the model successfully detected and segmented objects in the denoised images, which were completely missed in the raw inputs. While tested on a variety of images, some labels were incorrect since the YOLOv8 model had not been trained on the specific objects present in the scene. For example, traffic cones were occasionally misclassified as bottles, and a white storage cabinet was misclassified as a refrigerator. Similar misclassifications were also observed in the RGB stream, indicating that this is primarily a limitation of the YOLOv8 model rather than the proposed CLEAR-IR framework. Importantly, despite label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
IV-D ArUco Marker Detection
@@
Figure 8: Comparison of ArUco marker detection performance on RGB, raw IR, and denoised IR images.
@@
The CLEAR-IR framework addresses these challenges by enabling robust ArUco marker detection in IR images. By removing the spurious dot patterns, the model ensures the marker’s grid is clearly visible and interpretable by the detection algorithm, as demonstrated in Figure 8d. This is a significant advancement for low-light scenes where raw IR images are typically unusable for this task. The denoised output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping. This outcome directly contrasts with findings in prior work, such as [33], which concluded that raw IR images were unsuitable for low-light tag detection without the use of retroreflective materials. Similarly, this approach simplifies methods like [34], which propose using specialised tags for low-light localisation, by requiring no special tags or materials.
@@ V Conclusion
This work demonstrates that the CLEAR-IR framework effectively addresses a fundamental challenge in robotic perception by successfully removing structured light emitter patterns from infrared images. Leveraging a U-Net-based architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a zero-shot scenario. The significance of this contribution is further evidenced by the substantial improvement in performance for crucial downstream tasks, including feature matching, object detection, and marker detection. As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques. The CLEAR-IR framework thus provides a robust and reliable foundation for enabling autonomous systems to operate effectively in a wider range of challenging lighting conditions.
@@ VI Future Direction
For integration with vSLAM, future work will focus on assessing the effectiveness of CLEAR-IR within a complete visual SLAM pipeline. This will involve real-time deployment and a quantitative evaluation of its impact on trajectory estimation accuracy and loop closure detection in illumination-constrained environments.
@@ Acknowledgment
RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the main author with setup and data collection in a dark environment.

### v2

@@ IV Experimental Results
TABLE I: Performance comparison of image enhancement methods for various computer vision tasks.
Method Object Detection (Visible Objects) ArUco Marker Detection
Detected Missed Mislabelled Correct Missed
Retinex Theory [3] 1 9 0 5 1
CLAHE [27] 1 9 0 3 3
Zero-DCE [13] 0 10 0 5 1
LLFormer [29] 0 10 0 5 1
CIDNet [31] 1 9 0 6 0
Raw IR 0 9 1 0 6
U-Net 7 3 0 6 0
CLEAR-IR (Proposed Method) 11 0 2 6 0
Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively. Table I presents a comprehensive quantitative comparison of the performance of various image enhancement methods on high level tasks such as object detection, and ArUco marker detection. The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms. The proposed work is also used in feature tracking and odometry estimation to highlight the stability across frames in unseen environments.
IV-A Object Detection and Segmentation
@@
To demonstrate the utility of the reconstructed images for advanced perception tasks, a pre-trained YOLOv26 object detection and segmentation model was used for evaluation. This model was not trained on IR data, making this a zero-shot evaluation to determine if the proposed preprocessing step can make the images usable with standard RGB-based vision pipelines.
(a) YOLOv26 on RGB image
(b) YOLOv26 on raw IR image
(c) YOLOv26 on U-Net enhanced image
(d) YOLOv26 on CLEAR-IR image
@@
Figure 4: Object detection and segmentation results of YOLOv26 on raw and reconstructed IR images.
@@
On the raw IR images, YOLOv26 failed to detect or segment objects. The emitter patterns and resulting image noise prevented the model from recognising the distinct shapes and features of common objects. After preprocessing the images with the CLEAR-IR model, the performance of YOLOv26 dramatically improved. As illustrated in Figure 4, the model successfully detected and segmented objects in the reconstructed images, which were completely missed in the raw inputs. While tested on a variety of images, some labels were incorrect since the YOLOv26 model had not been trained on the specific objects present in the scene. For example, traffic cones were occasionally misclassified as bottles, and a white storage cabinet was misclassified as a refrigerator. Although the greyscale nature of IR data makes some misclassification inevitable without fine-tuning. Importantly, despite some label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
IV-B ArUco Marker Detection
@@
Figure 5: Comparison of ArUco marker detection performance on RGB, raw IR, and reconstructed IR images.
@@
The CLEAR-IR framework addresses these challenges by enabling robust ArUco marker detection in IR images. By removing the spurious dot patterns, the model ensures the marker’s grid is clearly visible and interpretable by the detection algorithm, as demonstrated in Figure 5d. This is a significant advancement for low-light scenes where raw IR images are typically unusable for this task. The reconstructed output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping. This outcome directly contrasts with findings in prior work, such as [9], which concluded that raw IR images were unsuitable for low-light tag detection without the use of retroreflective materials. Similarly, this approach simplifies methods like [25], which propose using specialised tags for low-light localisation, by requiring no special tags or materials.
IV-C Visual Simultaneous Localisation and Mapping (VSLAM) Performance
TABLE II: Quantitative evaluation of SLAM trajectory accuracy (RMSE in meters) and inference latency across Low Light (LL) and Extreme Low Light (ELL) scenarios.
Method Low Light (LL) - RMSE $\pm$ std Extreme Low Light (ELL) - RMSE $\pm$ std Inference (ms)
FB SQ1 SQ2 FB SQ1 SQ2 Mean std
[…]
Raw IR DNI DNI DNI DNI DNI DNI DNI DNI
[…]
DNI: Did Not Initialise; Bold shows best results
(a) SQ1 trajectories in the low-light sequence
(b) SQ1 trajectories in the extreme low-light sequence
Figure 6: Square 1 lap (SQ1) Trajectories
To evaluate the practical utility of CLEAR-IR for robotics applications, we integrated the reconstructed infrared stream into a modified ORB-SLAM3 [4] Visual SLAM pipeline. In our implementation, the default ORB feature extractor was replaced with SuperPoint [8], and feature matching was adapted accordingly to use floating-point descriptor matching with L2 distance. This modification ensures compatibility with learned feature descriptors while preserving the full SLAM pipeline, including tracking, local mapping, and loop closure. Robust SLAM performance critically depends on stable and repeatable feature tracking across frames. In raw IR images, however, the structured light emitter pattern introduces dense artificial high-frequency responses. These phantom features are inconsistent across viewpoints and frames, leading to incorrect correspondences, tracking divergence, and eventual tracking failure. This behaviour is reflected in the Table II where Raw IR fails in all sequences.
To ensure fair comparison with RGB-based enhancement methods, monocular RGB and depth sequences were recaptured under controlled lighting conditions. Three trajectories were recorded:
  • Forward-Backward (FB): This sequence is a simple movement of moving in a straight line till an object and reversing backward
  • Square 1 lap (SQ1): This sequence is a simple movement of rotating along an arena and returning to the start point as seen in Figure 6.
  • Square 2 lap (SQ2): This sequence repeats the SQ1 trajectory for two full laps, providing a benchmark for observing the model’s ability to facilitate successful loop closure and drift correction
Table II reports results in low-light conditions where limited ambient illumination is still present. Under these conditions, the RGB enhancement methods (Retinex, CLAHE, Zero-DCE, LLFormer, CIDNet) perform reliably because the RGB images retain sufficient structural information for feature extraction and tracking. CLEAR-IR and the U-Net baseline achieve performance comparable to these RGB-based methods. In simpler trajectories (FB and SQ1), CLEAR-IR performs on par with the RGB enhancement models. Notably, in the SQ2 sequence, CLEAR-IR achieves the lowest RMSE among all evaluated methods. Since SQ2 stresses loop closure and long-term consistency, this result indicates that CLEAR-IR reconstructs temporally stable structures that reduce drift accumulation over extended trajectories. Importantly, unlike raw IR, the reconstructed outputs do not introduce false high-frequency artefacts from the emitter pattern. Instead, the learned reconstruction produces geometrically meaningful and temporally consistent features suitable for SLAM.
Furthermore in extreme low-light conditions the RGB images become severely underexposed, containing insufficient texture or gradient information for reliable feature detection. Consequently, all RGB-based enhancement methods fail to maintain tracking across every sequence, as indicated by missing RMSE values. In contrast, the IR images remain unaffected by ambient lighting conditions. While raw IR still fails due to emitter induced phantom features, both U-Net and CLEAR-IR successfully enable full SLAM operation across all sequences. CLEAR-IR consistently outperforms the U-Net baseline in terms of RMSE, achieving the best accuracy in FB, SQ1, and SQ2. This demonstrates a key advantage of the proposed framework: when RGB information collapses, CLEAR-IR leverages the invariant structural information present in the IR modality to reconstruct geometrically consistent representations that remain usable for feature-based SLAM. In terms of inference latency, CLEAR-IR maintains competitive runtime performance. While classical methods such as CLAHE are significantly faster, they do not provide robustness under extreme lighting degradation. Conversely, transformer-based models such as LLFormer incur substantially higher computational cost. CLEAR-IR achieves a favourable trade-off between localisation accuracy and inference speed, maintaining compatibility with real-time SLAM update requirements.
@@ V Conclusion
Utilising the infrared cameras already commonly integrated into robotic systems, the CLEAR-IR framework effectively addresses a fundamental perception challenge by removing emitter patterns from infrared images, successfully enabling low-light vision and extending the use case of these camera systems. Leveraging a DeepMAO inspired architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a fully dark environment. The significance of this contribution is further evidenced by the substantial improvement in performance for downstream tasks, including VIO, object detection, and marker detection. As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques, showing flexibility even in entirely dark scenes. The CLEAR-IR framework thus provides a robust and reliable foundation for enabling autonomous systems to operate effectively in a wider range of challenging lighting conditions.
@@ Acknowledgment
RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the authors with setup and data collection in a dark environment.